OpenStudy (minimooe):

Please describe how you got the answer <3 Solve the system of the equations algebraically. y=x^2+2x y=3x+20

OpenStudy (razork):

y=3x+20 Step 1-Subtract 20 from both sides equally. ( y-20=3x) so it would be ( -20=3x) Step 2- Divide out your X's. (-20/3=3x/3) Thus, making it -6.66=X

OpenStudy (razork):

I'm pretty sure that's how ya do it. But I don't know about the top one.

OpenStudy (razork):

But since ur solving for Y, it could also be y=20+3x

rebeccaxhawaii (rebeccaxhawaii):

http://prntscr.com/b15xaw